### v1

# STORE: Streamlining Semantic Tokenization and Generative Recommendation with A Single LLM

Abstract.
Traditional recommendation models often rely on unique item identifiers (IDs) to distinguish between items, which can hinder their ability to effectively leverage item content information and generalize to long-tail or cold-start items. Recently, semantic tokenization has been proposed as a promising solution that aims to tokenize each item’s semantic representation into a sequence of discrete tokens. In this way, it preserves the item’s semantics within these tokens and ensures that semantically similar items are represented by similar tokens. These semantic tokens have become fundamental in training generative recommendation models. However, existing generative recommendation methods typically involve multiple sub-models for embedding, quantization, and recommendation, leading to an overly complex system. In this paper, we propose to streamline the semantic tokenization and generative recommendation process with a unified framework, dubbed STORE, which leverages a single large language model (LLM) for both tasks. Specifically, we formulate semantic tokenization as a text-to-token task and generative recommendation as a token-to-token task, supplemented by a token-to-text reconstruction task and a text-to-token auxiliary task. All these tasks are framed in a generative manner and trained using a single LLM backbone. Extensive experiments have been conducted to validate the effectiveness of our STORE framework across various recommendation tasks and datasets. We will release the source code and configurations for reproducible research.
Semantic Tokenization, Generative Recommendation, Large Recommendation Model
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
1. Introduction
Figure 1. Comparison between conventional unique identifiers (in light green) and semantic identifiers (in light purple). For simplicity, each semantic identifier in this illustration consists of only two tokens.
Recommender systems are widely deployed in online applications, such as e-commerce websites, advertising networks, streaming services, and social media, to deliver personalized recommendations tailored to users’ interests. However, traditional recommendation models often rely on unique item identifiers (IDs) to represent items (Yuan et al., 2023; Liu et al., 2024b), which presents several key limitations. First, ID-based item representation can suffer from overfitting due to the typically sparse and imbalanced nature of the training data. Second, it fails to fully leverage item content information, which is crucial for improving recommendations for long-tail and cold-start items. Third, ID embeddings are learned and updated independently, making it difficult to capture the semantic relationships among items. Finally, given the vast number of items, the item embedding table often consumes a significant amount of memory.
To address these limitations, semantic tokenization has recently emerged as a promising solution and has gained rapid traction in the community (Singh et al., 2024; Rajput et al., 2024; Liu et al., 2024a; Zhu et al., 2024). As illustrated in Figure 1, instead of representing each item with a unique ID embedding, semantic tokenization encodes each item’s semantic representation into a compact sequence of discrete tokens. These tokens, also known as semantic identifiers or codes [^1] In this paper, the terms semantic identifiers, discrete tokens, and codes may be used interchangeably., preserve the item’s semantics into the token space and ensure that semantically similar items are represented by similar tokens. For example, consider a scenario where each item is represented by a sequence of 4 consecutive tokens, each capable of assuming one of 256 possible values. This theoretically creates a representation space of approximately $256^{4}\approx 4$ billion combinations, which is sufficient or easily expandable for industrial-scale systems. Additionally, tokens and their embeddings can be shared across items, allowing the similarity between two items to be roughly estimated by the Hamming distance between their token sequences. This information can be obtained without the need for training on subsequent recommendation tasks and is independent of token embeddings. As a result, semantic tokenization effectively addresses the aforementioned challenges and shows promise in semantic-enhanced modeling for recommendations (Rajput et al., 2024; Singh et al., 2024).
With the recent success of generative AI, particularly large language models (LLMs) (Touvron et al., 2023; OpenAI, 2022), generative recommendation (Li et al., 2024; Rajput et al., 2024) has emerged as a new paradigm for item recommendation, framing it as a token sequence generation task. Semantic tokenization is a crucial preliminary step in training these models. Directly generating item ID sequences, as seen in models like SASRec (Kang and McAuley, 2018) and Bert4Rec (Sun et al., 2019), is challenging in practice due to the time-consuming process of decoding from the vast volume of item IDs. By tokenizing each item into a sequence of tokens, it becomes more efficient to generate over token sequences and subsequently map them back to items. This approach has significantly contributed to the success of recent generative recommendation models, such as TIGER (Rajput et al., 2024), EAGER (Wang et al., 2024b), and LC-REC (Zheng et al., 2024).
However, existing generative recommendation methods (e.g., TIGER (Rajput et al., 2024)) typically follow a pipeline that involves multiple distinct models: an embedder, a quantizer, and a recommender. As illustrated in the upper panel of Figure 2, the process begins with an embedder that extracts item semantic embeddings using a pre-trained text encoder. Next, a quantizer tokenizes these embeddings into discrete tokens, often using vector quantization techniques like RQ-VAE (Wang et al., 2024b; Singh et al., 2024). These tokens are then used to train the recommender through token sequence generation tasks, resulting in a complex, multi-stage pipeline with several inherent limitations: 1) Current approaches rely on frozen general-domain text encoders (e.g., Sentence-T5 (Ni et al., 2021)) for item embedding, which fail to address the domain gap between the general-domain corpus and item content in recommendation scenarios. 2) The commonly used RQ-VAE quantizer is challenging to train, requiring additional efforts to prevent codebook collapse and manage collisions (i.e., different items being assigned the same token sequence) (Rajput et al., 2024). 3) The pipeline involves the disjoint optimization of multiple models, leading to information loss and hindering knowledge transfer between them. 4) The complex nature of the pipeline can incur additional costs for model training, deployment, and management in practice.
In this paper, we introduce a unified framework, STORE, to streamline the processes of semantic tokenization and generative recommendation using a single LLM. Specifically, we formulate semantic tokenization as a text-to-token task and generative recommendation as a token-to-token task. To minimize information loss and enhance knowledge transfer, we incorporate a token-to-text reconstruction task for semantic tokenization and a text-to-token auxiliary task during generative recommendation. All tasks are framed in an instruction-based generative paradigm and trained on a domain-specific corpus using a single LLM backbone. Additionally, we utilize the classic non-parametric $k$-means algorithm for token discretization, simplifying the overall training process. Our trained model can be applied to various downstream recommendation tasks, such as sequential recommendation and click-through rate (CTR) prediction. In summary, our work makes the following main contributions:
Figure 2. Comparison of the existing semantic tokenization pipeline (embedder → quantizer → recommender) with our proposed pipeline (tokenizer → clusterer → recommender).
  • Unified Framework: We introduce STORE, a novel framework that streamlines semantic tokenization and generative recommendation using a single LLM. This approach minimizes information loss, enhances knowledge transfer, and reduces the pipeline complexity by using a single LLM model.
  • Efficient Semantic Tokenization: Our method employs a dense tokenizer to convert item content features into token embeddings, followed by k-means clustering to obtain discrete tokens. This design circumvents the challenges often encountered in training vector quantization models, such as codebook collapse (Huh et al., 2023; Baykal et al., 2023).
  • Empirical Evaluation: Extensive experiments on two real-world datasets – MIND for news recommendation and Yelp for restaurant recommendation – demonstrate that the STORE framework achieves superior performance across multiple recommendation scenarios, highlighting its effectiveness and broad applicability.
Figure 3. Detailed architecture for our designed dense tokenizer. Items are initially compressed into embeddings and subsequently clustered to generate semantic tokens, which are then incorporated into the downstream recommender system. Such “text-to-token” conversion will be achieved by cascaded attention mask mechanism and self-supervised tasks.
2. Preliminaries
Traditional recommender systems use unique item identifiers mapped to learnable vectors (embeddings) to represent items during training, leading to the challenges mentioned earlier. Recent studies (Rajput et al., 2024; Zheng et al., 2024; Liu et al., 2024a) have shown that semantic tokenization is becoming an increasingly popular approach for embedding item content into identifiers. [^2] Recent works (Wang et al., 2024b, a; Qu et al., 2024) have extracted item embeddings from pretrained recommenders using collaborative knowledge. However, these collaborative embeddings are unstable and frequently change in real-world scenarios, unlike the stable nature of item content knowledge. Therefore, this paper focuses on generating semantic identifiers merely based on item content features. In this section, we focus on two key aspects: i) generating semantic identifiers using standard approaches, and ii) utilizing these semantic identifiers in downstream recommendation models.
2.1. Semantic Tokenization
Semantic tokenization aims to generate shared sub-identifiers that can be linked across different items, based on item content or collaborative insights. As illustrated in Figure 2, the standard semantic tokenization pipeline (Rajput et al., 2024; Zheng et al., 2024) consists of an embedder, a quantizer, and a recommender.
The embedder, typically a pretrained language model such as SentenceBERT (Reimers and Gurevych, 2019) or LLaMA (Touvron et al., 2023), is often used because textual features are the primary modality in most recommendation scenarios. Leveraging the rich knowledge from pretrained corpora, these embedders generate robust item representations.
Next, a residual quantizer (Lee et al., 2022) is used to discretize each item embedding into structure-aware tokens, using multi-layer ($L$) codebooks, with each codebook containing $K$ code vectors. Each item is (expected to be) mapped to a combination of codes with a length of $L$, where each position is selected from the corresponding codebook. Theoretically, the representation space for $L$-layer codebooks is $K^{L}$, which means that even a much smaller $K$ and $L$ can effectively represent a total of $N$ items, even when $N\gg K$. Consequently, the substantial memory required for item embeddings in traditional recommenders, i.e., $N\times D$ where $D$ is the embedding dimension, can be compressed into a significantly smaller, logarithmic space, i.e., $K\times L\times D$.
2.2. Downstream Recommenders
Once the semantic codes are obtained, they can be integrated into various recommendation models. In sequential recommendation (Rajput et al., 2024; Wang et al., 2024b), the user sequence is replaced by a flattened sequence of semantic codes, shifting the task from next-item prediction to next-code prediction. Alternatively, these codes can enhance the recommendation ability of large language models by integrating collaborative knowledge (Zheng et al., 2024; Qu et al., 2024).
3. Proposed Approach: STORE
In this section, we introduce our STORE model, which includes a dense tokenizer, a simple clusterer, and a recommender. This design addresses the limitations of the standard paradigm.
Figure 4. Cascaded attention mask for dense tokenizer. “Ph. Block” represents the placeholder block.
Firstly, our dense tokenizer compresses item content into several embeddings, enabling the reconstruction of the original content with minimal information loss. This is achieved by post-training a large language model on domain-specific data, which also reduces the distribution gap between the language model and the recommendation scenario. Secondly, simple, training-free algorithms can be applied to cluster the positions of multiple dense embeddings. Unlike residual quantizers, which may suffer from codebook collapse, this simplicity ensures balanced distribution across clusters and a high utilization rate. Thirdly, the downstream recommender shares the same backbone – a large language model – as the tokenizer, since the semantic tokens are generated based on the model’s textual comprehension. Therefore, item semantic knowledge will be transferred not only explicitly through the semantic tokens but also implicitly through the shared network parameters.
3.1. Dense Tokenizer
To distill content into short tokens, we introduce a dense tokenizer compatible with any decoder-only large language model. This distillation leverages a block-wise input scheme, hierarchical attention masking scheme, and self-supervised post-pretraining tasks. The design of the dense tokenizer is inspired by the gisting framework (Mu et al., 2024), which compresses prompts into short tokens to enhance inference efficiency in the natural language processing domain.
Figure 5. Instruction tuning templates.
Block-wise Input Scheme. The input sequence is organized into four distinct blocks: content, token, placeholder, and task. Given an item $\mathbf{x}$ – for example, a news article – with $m$ attributes $\{\mathbf{a}_{1},\mathbf{a}_{2},\cdots,\mathbf{a}_{m}\}$, where $\mathbf{a}_{i}$ represents the $i$-th attribute (e.g., title, abstract, category), the content block for the dense tokenizer is constructed as:
[…]
where [;][;][ ; ] denotes concatenation and $r\leq m$. An example content block might be “title: Yellowstone tourist injured ... abstract: A tourist suffered severe burns ...”.
Next, the token block contains $v$ fixed tokens:
[…]
where $v=2$ in Figure 3. These tokens have learnable embeddings that absorb knowledge through a multi-layer transformer network, serving as a bridge between the original content and the task output.
Since the output embeddings of the token block are expected to encapsulate the knowledge of the item content, they must also reconstruct or generate the original item attributes. To achieve this, we append a placeholder block containing $v$ placeholder tokens after the token block:
[…]
Unlike other blocks, which have embedding tables to map tokens into embeddings as the final input to the large language model, the placeholder block is filled with the corresponding output embeddings of the token block. This ensures that the task block is guided to generate text from the first transformer layer based on these output embeddings.
Finally, the input sequence ends with the task block, which includes a task token and the answer sequence:
[…]
where $t_{i}$ is a special token representing the $i$-th self-supervised task: it is a reconstruction task when $0<i\leq r$ and a generation task when $r<i\leq m$. For example, this could be: ‘‘reconstruct_title: title: Yellowstone tourist injured ...’’
Cascaded Attention Mask. Decoder-only large language models typically employ a causal attention mask, ensuring that each token in a sequence can only attend to preceding tokens and itself, but not to future tokens. This conventional approach is unsuitable for our scenario where only the outputs of the token block (and subsequently the placeholder block) are permitted to generate the task output. Therefore, we introduce a cascaded attention masking scheme that includes both inner-block and inter-block masking.
As illustrated in the diagonal of Figure 4, inner-block masking consistently enforces causal attention to preserve sequential knowledge comprehension. Conversely, inter-block masking can be configured as either full or empty attention: the content block fully attends to the token block, and the placeholder block fully attends to the task block. Attention between other blocks is prohibited and set to empty.
Post-pretraining. As depicted in Figure 3, we replicate each item content $m$ times to create $m$ training samples. Each sample is tailored to a specific reconstruction or generation task $t_{i}$.
We will employ low-rank adaptation (LoRA) (Hu et al., 2021), a parameter-efficient fine-tuning approach, to train the language model. Additionally, we will freeze the pretrained word embeddings while tuning the dense token embeddings and task embeddings.
To perform the filling operation on the placeholder block, we will conduct dual forward propagation: first to capture the output from the token block to fill the placeholder, and then to tune the language model with the next token prediction task:
(5) $\hat{a}_{i,j+1}=\text{argmax}_{w\in\mathbf{W}}P(w|a_{i,1},a_{i,2},\dots,a_{i,j}),$
optimized using cross-entropy loss, where $\mathbf{W}$ denotes the token vocabulary and $a_{i,j}$ denotes the $j$-th token of the attribute $\mathbf{a}_{i}$.
Notably, our approach differs the gisting framework in several ways: i) our compression targets item content in recommendation scenarios; ii) we introduce a placeholder block adjacent to the token block to ensure the task block is guided by the output of the token block; iii) we design both reconstruction and generation tasks based on the dense tokens.
Figure 6. Code tree used in conditional beam search.
3.2. Simple Clusterer
Unlike standard pipeline that employ complex, hard-to-train differentiable vector quantization techniques to segment item content embeddings into discrete tokens, our dense tokenizer efficiently maps item content features into reconstructable embeddings–termed dense tokens–that encapsulate domain-specific content. Hence, these dense vectors can be discretized into cluster indices using a training-free clustering approach. We first aggregate the output of the dense tokens for all items as follows:
(6) $\mathbf{E}=\begin{bmatrix}\mathbf{{\color[rgb]{0.5859375,0.44140625,0.21484375}\definecolor[named]{pgfstrokecolor}{rgb}{0.5859375,0.44140625,0.21484375}e}}_{1,1}&\mathbf{{\color[rgb]{0.5859375,0.44140625,0.21484375}\definecolor[named]{pgfstrokecolor}{rgb}{0.5859375,0.44140625,0.21484375}e}}_{1,2}&\cdots&\mathbf{{\color[rgb]{0.5859375,0.44140625,0.21484375}\definecolor[named]{pgfstrokecolor}{rgb}{0.5859375,0.44140625,0.21484375}e}}_{1,n}\\ \mathbf{{\color[rgb]{0.1796875,0.36328125,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1796875,0.36328125,0.52734375}e}}_{2,1}&\mathbf{{\color[rgb]{0.1796875,0.36328125,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1796875,0.36328125,0.52734375}e}}_{2,2}&\cdots&\mathbf{{\color[rgb]{0.5859375,0.44140625,0.21484375}\definecolor[named]{pgfstrokecolor}{rgb}{0.5859375,0.44140625,0.21484375}e}}_{2,n}\\ \vdots&\vdots&\ddots&\vdots\\ \mathbf{e}_{v,1}&\mathbf{e}_{v,2}&\cdots&\mathbf{e}_{v,n}\end{bmatrix},$
where $n$ denotes the number of items, and $\mathbf{e}_{i,j}$ represents the $i$-th output embedding of the dense tokens of the $j$-th item with $D$ dimensions.
Next, we apply the principal component analysis (PCA) technique (Maćkiewicz and Ratajczak, 1993) to reduce the high-dimensional $D$ vectors (over 1024 dimensions) to a lower dimension $d$, such as 32. The principal components are denoted by $\mathbf{\hat{e}}_{i,j}$ for each reduced embedding of $\mathbf{e}_{i,j}$.
Finally, we apply a simple, training-free clustering algorithm, such as K-Means (Krishna and Murty, 1999), to each row of the matrix, denoted as $\mathbf{\hat{E}}[i,:]=[\mathbf{\hat{e}}_{i,1},\mathbf{\hat{e}}_{i,2},\cdots,\mathbf{\hat{e}}_{i,n}]$. The resulting cluster indices are organized into a matrix:
(7) $\mathbf{C}=\begin{bmatrix}c_{1,1}&c_{1,2}&\cdots&c_{1,n}\\ c_{2,1}&c_{2,2}&\cdots&c_{2,n}\\ \vdots&\vdots&\ddots&\vdots\\ c_{v,1}&c_{v,2}&\cdots&c_{v,n}\end{bmatrix},$
where $1\leq c_{i,j}\leq k$ represents the cluster indices, and $k$ is the number of clusters. Thus, each item can be represented by $v$ discrete tokens: $\mathbf{c}_{j}=[c_{1,j},c_{2,j},\ldots,c_{v,j}]$.
3.3. Generative Recommender
Like other item tokenization approaches, the discrete tokens generated by the clusterer can be utilized in generative sequential recommendation. Since the dense tokens are derived from a large language model, we employ the same model as the backbone for both scenarios. Additionally, we have designed a text-token alignment task to enhance the token comprehension capabilities of large language models.
Training. We construct a new vocabulary incorporating the semantic tokens. As illustrated in Figure 5, the training involves the instruction tuning for both sequential recommendation and an additional text-token alignment tasks. The backbone model is also equipped with low-rank adaptation technique.
Inference. We devise conditional beam search to generate the next item. Initially, we construct a code (i.e., token) tree, as illustrated in Figure 6, which accommodates all possible code combinations. During the beam search, we set the code logits from the classification module to zero for any combinations that do not appear in the tree paths.
4. Experiments
4.1. Experimental Setup
Table 1. Dataset statistics.
MIND Yelp
#Items 25,634 73,380
#Users 45,000 45,000
#Finetune 40,000 40,000
#Test 5,000 5,000
Avg. User Length 11.78 6.47
Avg. Item Appearance 20.69 3.97
Datasets. We conduct experiments on two real-world content-based recommendation dataset, i.e., news recommendation dataset MIND (Wu et al., 2020a) and restaurant recommendation dataset Yelp [^3] https://www.yelp.com/dataset. The dataset statistics are summarized in Table 1.
Baseline models. We benchmark our proposed STORE framework against unique id-based recommenders (SASRec (Kang and McAuley, 2018) and P5 (Geng et al., 2022)) and semantic code-based recommenders (TIGER (Rajput et al., 2024) and LC-Rec (Zheng et al., 2024)). Specifically, SASRec employs a multi-layer attention network with causal attention, utilizing unique identifiers and trained on a next-item prediction task. P5 integrates unique item identifiers into large language models. TIGER and LC-Rec use pretrained SentenceBERT (Reimers and Gurevych, 2019) and Llama-1 to extract item content embeddings, respectively. For a fair comparison, all code-based recommenders, i.e., TIGER, LC-Rec, and our STORE, represent each item using four codes, with a fixed code vocabulary of 256 at each position.
Evaluation Metrics. We follow the common practice (Rajput et al., 2024; Qu et al., 2024) to evaluate the effectiveness of sequential recommenders with the widely used metrics, i.e., Recall and NDCG (Järvelin and Kekäläinen, 2002). In this work, we use Recall@1, Recall@5, Recall@10, Recall@20, NDCG@1, NDCG@5, NDCG@10, and NDCG@20 for evaluation.
Implementation Details. i) Tokenizer. We utilize the pretrained OPT-base (Zhang et al., 2022) as the backbone large language model for dense tokenization. Optimization is performed using the Adam (Kingma and Ba, 2015) optimizer with a learning rate of 1e-4, a batch size of 512, a LoRA rank of 32, and a token block size $v$ of 4. Self-supervised tasks are detailed in Table 2. ii) Clusterer. We apply PCA (Maćkiewicz and Ratajczak, 1993) to reduce 1024-dimensional item embeddings to 32 components, subsequently clustering each position into 256 groups. iii) Recommender. We set the maximum length of user history sequence to 20 and we use the last item in the sequence as the prediction target. We use the same pretrained OPT-base as the backbone with a learning rate of 5e-4, a batch size of 64 and a LoRA rank of 128. The training starts with the joint learning of generative recommendation task and text-token alignment task. After model convergence, the model will further be tuned by the single generative recommendation task. Early stopping mechanism is used with patience of 5. All the experiments are conducted on a single NVIDIA A100 device with 80GB memory. We release all our code and data here [^4] https://anonymous.4open.science/r/STORE/ for other researches to reproduce our work. We employ Recability [^5] https://github.com/Recability, a benchmark for evaluating the recommendation abilities of large language models, for most of our experiments.
Table 2. Statistics of post-pretraining tasks.
Content Block Generation Target
MIND ($m$=4, $k$=2) (title, abstract) title
(title, abstract) abstract
(title, abstract) category
(title, abstract) subcategory
Yelp ($m$=4, $k$=3) (name, city, address) name
(name, city, address) city
(name, city, address) address
(name, city, address) state
Table 3. Overall performance comparison in retrieval scenarios. We use R and N to represent the Recall and NDCG metrics, respectively.
MIND Yelp
Identifier Backbone R@5 R@10 R@20 N@5 N@10 N@20 R@5 R@10 R@20 N@5 N@10 N@20
Unique ID SASRec 0.0076 0.0086 0.0096 0.0181 0.0242 0.0308 0.0000 0.0002 0.0002 0.0002 0.0002 0.0003
P5 0.0068 0.0082 0.0089 0.0158 0.0204 0.0249 0.0000 0.0002 0.0002 0.0002 0.0002 0.0003
TIGER Transformer 0.0168 0.0256 0.0370 0.0422 0.0565 0.0717 0.0044 0.0064 0.0082 0.0113 0.0152 0.0193
TIGER OPT-base 0.0173 0.0272 0.0394 0.0424 0.0577 0.0720 0.0044 0.0062 0.0088 0.0114 0.0150 0.0198
LC-Rec OPT-base 0.0288 0.0428 0.0594 0.0654 0.0875 0.1111 0.0082 0.0108 0.0140 0.0192 0.0257 0.0327
STORE (ours) OPT-base 0.0726 0.0746 0.0764 0.1785 0.2389 0.3033 0.0224 0.0266 0.0316 0.0586 0.0785 0.0996
4.2. Generative Recommenders for Retrieval
Table 3 provides an overview of the performance across four baselines over two datasets. Drawing from the results, we can derive the following observations:
Firstly, the semantic code group (i.e., TIGER, LC-Rec, and STORE) consistently outperforms the unique identifier group (i.e., SASRec and P5). This superiority stems from several factors: i) Learning unique identifiers for items depends heavily on rich interaction signals, whereas semantic codes can be learned from multiple items sharing common codes, enhancing learning efficiency. ii) Semantic codes incorporate content features into sequential recommenders, which is not the case with unique identifiers.
Secondly, the Yelp dataset exhibits worse performance compared to the MIND dataset, as indicated by shorter user sequence lengths and a higher number of distinct items, which results in a lower average item appearance as shown in Table 1. Additionally, using identical settings (4 tokens, each with a vocabulary size of 256) for tokenization, the MIND dataset, with fewer items, is more readily distinguishable. Furthermore, the item content in the MIND dataset, which includes news titles and abstracts, is more informative than that in the Yelp dataset, which consists of restaurant names, cities, and addresses.
Thirdly, compared to the Transformer backbone, the OPT-based backbone achieves better performance using the same TIGER code, due to its superior deep context comprehension ability.
Fourthly, despite Llama’s larger size and its superior performance over the OPT-base model in various NLP tasks, our STORE exceeds other baselines, including LC-Rec, which employs Llama for item embedding extraction. This underscores the effectiveness of STORE and the benefits of our post-pretraining strategy on dense tokens.
Figure 7. Instruction tuning for scoring scenarios.
4.3. Ablation Study
Here, we study the effectiveness of various components within our framework. Based on the results from Table 4, we can make the following observations:
Firstly, STOREw/o dense ${}_{\text{w/o dense tokenizer}}$ follows the standard semantic tokenization pipeline: utilizing a pretrained OPT-base model to derive a single content embedding per item, which is then discretized into short tokens via RQ-VAE. Subsequently, these tokens are used to train the OPT-base model on a generative retrieval task. Our STORE surpasses this variant, demonstrating the superiority of our proposed paradigm.
Secondly, STOREw/o conditional beam ${}_{\text{w/o conditional beam search}}$ employs soft beam search constraints as used by LC-Rec (Zheng et al., 2024), where the classification at each position during next-code prediction is confined to the codebook size. While these soft constraints do enhance search performance, they cannot prevent the generation of token combinations that do not correspond to an actual item. For example, 2-1 shown in Figure 6 is not a valid path (token combination), yet it can be generated as valid under soft constraints during the generation process. In contrast, our STORE achieves a significant improvement, thereby validating the effectiveness of conditional beam search.
Thirdly, STOREw/o text-token alignment ${}_{\text{w/o text-token alignment task}}$, trained solely with the generative retrieval task (i.e., next-token prediction), performs less effectively compared to our STORE. This discrepancy likely arises because, despite using the same language model for both tokenization and recommendation, the model struggles to recognize randomly-initialized token embeddings. Introducing the text-token alignment task is crucial for bridging the gap between text and tokens, thereby enhancing the model’s semantic understanding of user sequences.
4.4. Generative Recommenders for Scoring
To illustrate the versatility of semantic codes, we examine their performance in a popular scenario using a large language model as a recommender: score prediction. As depicted in Figure 7, the language model processes a user sequence and a candidate item, predicting the likelihood of a user clicking on the item. The logits for the “yes” and “no” tokens at the final output position are selected, and then normalized with a softmax function. The final score is represented by the “yes” score. [^6] For GPT-3.5, we assign a score of 1 for a response of “YES” and 0 for “NO”.
Results for three groups are presented in Table 5. Zero-shot LLM recommenders such as BERT (Devlin, 2018), OPT (Zhang et al., 2022), Llama (Touvron et al., 2023), and GPT-3.5 (OpenAI, 2022) are not fine-tuned on the recommendation dataset. However, finetuning with textual features is computationally expensive: lengthy item content leads to extended user sequences, and computational costs increase quadratically with the length of the input sequence. By utilizing semantic codes, large language models become more tunable: if the average item content length is 40 and the semantic code length is 4, this reduction of 10 times in length theoretically results in a 100-fold acceleration of the attention module. Based on the results, we have the following observations:
Firstly, zero-shot large language models exhibit the poorest performance among the three groups. Despite their robust textual comprehension abilities, they lack exposure to collaborative corpora during pretraining, which hampers their recommendation ability. Notably, even advanced models like Llama and GPT-3.5, with their deeper networks and higher dimensionality, perform comparably to the zero-shot BERT-base model.
Secondly, the fine-tuned LLM recommender, such as BERT-base, achieves over a 20% improvement compared to its zero-shot counterpart. This highlights the necessity of fine-tuning language models with recommendation tasks to enhance their efficacy as recommenders. Due to BERT’s limited maximum sequence length, we truncate the user sequence, making fine-tuning somewhat feasible given its relatively small network size. However, attempting to finetune larger language models in a similar manner is unacceptable as mentioned above.
Thirdly, our generative recommender surpasses the performance of the other two groups. Although the semantic codes are new tokens for the language models, their advanced contextual comprehension capabilities enable the capture of both collaborative signals and content knowledge embedded within these codes. Crucially, replacing long item content with short semantic tokens significantly reduces computational demands, making the finetuning possible on larger language models like OPT-base, as opposed to BERT-base.
Fourthly, our STORE surpasses TIGER and LC-Rec in the scoring scenario, underscoring the high quality of codes generated in our proposed paradigm.
Table 4. Ablation studies. Experiments are conducted on the MIND dataset.
R@5 R@10 R@20 N@5 N@10 N@20
STORE w/o dense ${}_{\text{w/o dense tokenizer}}$ 0.0664 0.0670 0.0676 0.1685 0.2255 0.2863
STORE w/o conditional beam ${}_{\text{w/o conditional beam search}}$ 0.0296 0.0422 0.0602 0.0662 0.0868 0.1128
STORE w/o text-token alignment ${}_{\text{w/o text-token alignment task}}$ 0.0705 0.0722 0.0735 0.1758 0.2340 0.2592
STORE 0.0726 0.0746 0.0764 0.1785 0.2389 0.3033
Table 5. Overall performance comparison in scoring scenarios. We use R and N to represent the Recall and NDCG metrics, respectively.
Identifier Backbone AUC MRR N@1 N@5
Zero-shot LLM Recommender Text BERT-base 0.4963 0.4139 0.2655 0.3729
Text OPT-base 0.5490 0.4658 0.3715 0.4362
Text Llama-1 0.4583 0.3858 0.2100 0.3301
Text GPT-3.5 0.5057 - - -
Fine-tuned LLM Recommender Text BERT-base 0.6014 0.5055 0.4178 0.4890
Generative Recommender TIGER OPT-base 0.6202 0.5277 0.4987 0.5315
LC-Rec OPT-base 0.6043 0.5052 0.4543 0.4988
STORE (ours) OPT-base 0.6505 0.5509 0.5062 0.5542
5. Related Work
5.1. LLMs for Recommendation
Generally, the emerging techniques of LLMs for enhancing recommender systems can be grouped into three paradigms, namely pre-training, prompting, and fine-tuning (Zhao et al., 2024).
Pre-training.
Research in this paradigm typically involves tasks designed to model diverse user behaviors and aims to develop a fundamental recommendation model. For instance, PTUM (Wu et al., 2020b) employs two pre-training tasks: masked behavior prediction and next K behavior prediction. Similarly, Cui et al. (Cui et al., 2022) introduce M6, which utilizes an auto-regressive generation task and a text-infilling objective. Additionally, Geng et al. (Geng et al., 2022) propose P5, a model that integrates multiple recommendation tasks within a unified framework to pre-train a foundational recommendation model.
Prompting.
Instead of pre-training an LLM, some studies aim to directly integrate LLMs into the recommendation pipeline without parameter updates, typically through feature augmentation. For instance, Xi et al. (Xi et al., 2023) propose utilizing LLMs to infer user preferences and factual knowledge about items. Similarly, Wang et al. (Wang and Lim, 2023) employ LLMs to model user preferences.
Fine-tuning.
This line of research seeks to leverage the capabilities of existing powerful LLMs with fine-tuning. Fine-tuning is a critical step in aligning LLMs with various downstream recommendation tasks. Related studies either adopt full-model fine-tuning (Friedman et al., 2023; Shen et al., 2023) or employ parameter-efficient fine-tuning techniques (Bao et al., 2023; Wu et al., 2024), such as LoRA (Hu et al., 2021), to reduce computational resource requirements.
5.2. Generative Recommendation
Generative recommenders learn from user interactions or sequential patterns to directly generate recommendations without the need for filtering or ranking (Li et al., 2023). Traditional approaches such as SASRec (Kang and McAuley, 2018) and BERT4Rec (Sun et al., 2019), along with language model-based recommenders like P5 (Geng et al., 2022) and VIP5 (Geng et al., 2023), utilize unique identifiers to represent items and generate the next item by selecting the most probable item from the entire distribution.
To incorporate item content knowledge, TIGER (Rajput et al., 2024) introduced semantic identifiers that can be shared across different items, replacing the unique identifier. This approach has been further developed by other semantic tokenization efforts (Zheng et al., 2024; Jin et al., 2023; Liu et al., 2024a). Moreover, the effectiveness of integrating collaborative features learned from simple recommenders into identifiers has been demonstrated (Qu et al., 2024; Wang et al., 2024a). However, this method heavily relies on rich interactions and tends to be unstable or change frequently in real-world scenarios. The learned tokens are used in generative recommenders, shifting the training focus from next-item prediction to next-code prediction. This shift narrows the search space for each position, thereby enhancing the inference performance. Given the limitations discussed in the Preliminaries Section, this paper aims to reevaluate and refine the standard semantic tokenization pipeline.
6. Conclusion
In this paper, we introduce the STORE framework, which streamlines semantic tokenization and generative recommendation using a single LLM. Unlike existing methods that rely on separate sub-models for embedding, quantization, and generation, our STORE framework unifies these tasks within a single generation framework, simplifying the overall process by reusing the LLM backbone. Our experimental results demonstrate that the STORE framework outperforms existing baselines across both retrieval and scoring tasks on real-world recommendation datasets. Additionally, the versatility of the STORE framework extends beyond purely text-based applications, showing promise in multimodal domains.